iosangel:

What is the perimeter of a polygon with vertices at (-3, 1), (5, 1), (-3, 4), (5, 4

darkknight:

try making a drawing, and then it will be easier to find the perimeter, you can graph and it would help

darkknight:

where ya at now?

iosangel:

im graphing

darkknight:

alright, keep working on it.

iosangel:

i got 16.54 but i think its wrong

darkknight:

How about you just take the distance from each point to the next and add them up? see what u get

darkknight:

Yeah you shouldn't get a decimal

iosangel:

Thats what i did i used the distance formula to find the distance between the points and thats where i think i got something wrong, here let me show you my work

XioGonz:

Do you have a photo of any sort?

iosangel:

AB=√(5--3)^2-(1-1)^2=√64=8 BC=√(-3-5)^2-(4-1)^2=√-73=8.54 cd would be the same as AD and ad would be the same as BC Then i added 8.54 and 8 and got 16.54

iosangel:

i mean CD would be the same as AB

XioGonz:

I got 21 but I might not be doing this with the same formula you are.

iosangel:

The distance formula √(x2-x1)^2-(y2-y1)^2

iosangel:

I dont know what im doing wrong

darkknight:

you don't need a distance formula. This is a lot simpler of a problem than you think. If you look at the following coordinates, some match with their x-values and some- with y values. (-3, 1), (5, 1), (-3, 4), (5, 4 The distance between (-3,1) and (5,1) is 8. And the distance between (-3, 1 and -3,4) is 3. These are 2 sides of our rectangle. now use 2(length) + 2(width) = perimeter to solve for the perimeter

iosangel:

oh so it would be 22

darkknight:

yes.

iosangel:

thank you, your way helps

darkknight:

np.